OpenStudy (anonymous):

Find the distance between these coordinates... (-1,-3); (11,7); d=???

OpenStudy (uri):

Distance formula.

OpenStudy (uri):

Do you know?

OpenStudy (anonymous):

i do not

OpenStudy (uri):

(x1,y1)=-1,-3 (x2,y2)=11,7 Put the values in the formula and try doing it.

OpenStudy (uri):

@countonme123

OpenStudy (anonymous):

i did i got square root of 244? is that right???

OpenStudy (uri):

Let me do it now.

OpenStudy (uri):

Do you have options?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok let me know and maybe you can explain where i screwed up

OpenStudy (uri):

Yes under root 244 is right :D

OpenStudy (phi):

(-1,-3); (11,7); -1 -11 = -12 , (-12)^2 = 144 -3 -7 = -10 (-10)^2 =100 add and get 244 sqrt(244) is the answer. Looks like you did it correctly

OpenStudy (anonymous):

hmm but i got it wrong when i put that. is there a way to simplify that

OpenStudy (uri):

@phi o_O you're coool,no? *heavy sarcasm*

OpenStudy (phi):

right away we can write 244 as 2*122 = 2*2*61 so 2 sqrt61)

OpenStudy (anonymous):

you mean 4?

OpenStudy (phi):

sqrt(4*61) = 2 * sqrt(61)

OpenStudy (uri):